# Strong convergence of fully discrete finite element schemes for the stochastic semilinear generalized Benjamin-Bona-Mahony equation driven by additive Wiener noise

### v1

@@ 4. Proof of Main Results
$\begin{split}\mathbb{E}\|u_{h}(t)\|^{2}&=\mathbb{E}\|(\mathcal{P}_{h}+A_{h})^{-1}v_{h}(t)\|^{2}\\ &\leq\mathbb{E}\|v_{h}(t)\|^{2}\\ &\leq B(t)+3Ct\int_{0}^{t}\mathbb{E}\|u_{h}(\tau)\|^{2}d\tau\end{split}$
@@
Now we prove our first main result of the paper. First, we find the error between $v(t)$ and $v_{h}(t)$. Using this error, we obtain the rate of convergence of finite element approximation of stochastic semilinear generalized BBM equation (2.1).
@@
$I_{1}(t)\leq Ch^{\beta}\left(\|v_{0}\|_{L_{2}(\Omega,\dot{H}^{\beta})}+\|A^{(\beta-1)/2}Q^{1/2}\|\right).$ (4.3)
@@
$\begin{split}I_{2}(t)&=\left\|\int_{0}^{t}\left(E(t-\tau)f(u(\tau))-E_{h}(t-\tau)\mathcal{P}_{h}f(u_{h}(\tau))\right)d\tau\right\|_{L_{2}(\Omega,H)}\\ &=\bigg{\|}\int_{0}^{t}\bigg{[}E(t-\tau)f(u(\tau))-E_{h}(t-\tau)\mathcal{P}_{h}f(u(\tau))\\ &\hskip 56.9055pt+E_{h}(t-\tau)\mathcal{P}_{h}f(u(\tau))-E_{h}(t-\tau)\mathcal{P}_{h}f(u_{h}(\tau))\bigg{]}d\tau\bigg{\|}_{L_{2}(\Omega,H)}\\ &\leq\int_{0}^{t}\left\|\left(E(t-\tau)-E_{h}(t-\tau)\mathcal{P}_{h}\right)f(u(\tau))\right\|_{L_{2}(\Omega,H)}d\tau\\ &+\int_{0}^{t}\left\|E_{h}(t-\tau)\mathcal{P}_{h}(f(u(\tau))-f(u_{h}(\tau)))\right\|_{L_{2}(\Omega,H)}d\tau\\ &\leq Ch^{\beta}\int_{0}^{t}(t-\tau)^{-\beta/2}\|f(u(\tau))\|_{L_{2}(\Omega,H)}d\tau\\ &+\int_{0}^{t}\|f(u(\tau))-f(u_{h}(\tau))\|_{L_{2}(\Omega,H)}d\tau\\ &\leq Ch^{\beta}\int_{0}^{t}(t-\tau)^{-\beta/2}(1+\|u(\tau)\|_{L_{2}(\Omega,H)})d\tau\\ &+C\int_{0}^{t}\|u(\tau)-u_{h}(\tau)\|_{L_{2}(\Omega,H)}d\tau\\ &\leq Ch^{\beta}+C\int_{0}^{t}\|u(\tau)-u_{h}(\tau)\|_{L_{2}(\Omega,H)}d\tau.\end{split}$
@@
To estimate $J(t)$, we use the connection between the resolvent of the generator and the corresponding $C_{0}$-semigroup [13, Theorem 4.2]. For $t\geq 0$, we get
@@
This completes the proof of Theorem (2.3). ∎
@@
The solution of the stochastic parabolic component $v(t)$ is regular in time ( see [21, Lemma 2.7] and [15, Proposition 3.4]) i.e.
@@
$\begin{split}J_{2}&=\sum_{j=0}^{n-1}\int_{t_{j}}^{j+1}\|R(kA_{h})^{n-j}\mathcal{P}_{h}f(U^{j})\|_{L_{2}(\Omega,H)}d\tau\\ &=\sum_{j=0}^{n-1}\int_{t_{j}}^{j+1}\left(\mathbb{E}\|R(kA_{h})^{n-j}\mathcal{P}_{h}f(U^{j})\|^{2}\right)^{1/2}d\tau\\ &\leq K_{f}C\sum_{j=0}^{n-1}\int_{t_{j}}^{j+1}\left(\mathbb{E}\|U^{j}\|^{2}\right)^{1/2}d\tau\\ &=K_{f}Ck\sum_{j=0}^{n-1}\left(\mathbb{E}\|U^{j}\|^{2}\right)^{1/2}\\ &\leq K_{f}Ck\sqrt{n}\left(\sum_{j=0}^{n-1}\mathbb{E}\|U^{j}\|^{2}\right)^{1/2}.\end{split}$
@@
$\|A^{\nu}E(t)\|\leq Ct^{-\nu},\quad\text{ for }t>0$ (4.5)
$\|A^{\mu}(I-E(t))\|\leq Ct^{\mu},\quad\text{ for }t\geq 0.$ (4.6)
Now we prove our second main result. First, we will find the full discretization error between $V^{n}$ and $v(t_{n})$ for any $t_{n}$ in $[0,T]$. By using this error, we will find the full discretization error for the stochastic semilinear generalized BBM-Burger equation.
@@
$\begin{split}Err_{d,1}&=\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}(\mathbb{E}\|R(kA_{h})^{n-j}\mathcal{P}_{h}(f(U^{j})-f(u(t_{j})))\|^{2})^{1/2}d\tau\\ &\leq Ck\sum_{j=0}^{n-1}(\mathbb{E}\|f(U^{j})-f(u(t_{j}))\|^{2})^{1/2}\\ &\leq CK_{f}k\sum_{j=0}^{n-1}(\mathbb{E}\|U^{j}-u(t_{j})\|^{2})^{1/2}\\ &\leq C\sqrt{n}k\left(\sum_{j=0}^{n-1}\mathbb{E}\|U^{j}-u(t_{j})\|^{2}\right)^{1/2}.\end{split}$
@@
$\begin{split}Err_{d,2}&=\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}(\mathbb{E}\|R(kA_{h})^{n-j}\mathcal{P}_{h}(f(u(t_{j}))-f(u(\tau)))\|^{2})^{1/2}d\tau\\ &\leq CK_{f}\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}(\mathbb{E}\|u(t_{j})-u(\tau)\|^{2})^{1/2}d\tau\\ &\leq CK_{f}\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}(t_{j}-\tau)^{\gamma/2}d\tau\\ &\leq Ck^{\gamma/2}\end{split}$
For the estimate of $Err_{d,3}$, we apply Lemma $2.8$ of [21] and (ii) to obtain
$\begin{split}Err_{d,3}&=\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}(\mathbb{E}\|R(kA_{h})^{n-j}\mathcal{P}_{h}-E(t_{n}-t_{j})f(u(\tau))\|^{2})^{1/2}d\tau\\ &=\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}(\mathbb{E}\|F_{n-j}A^{(1-\beta)/2}A^{(\beta-1)/2}f(u(\tau))\|^{2})^{1/2}d\tau\\ &\leq\sum_{j=0}^{n-1}\|F_{n-j}A^{(1-\beta)/2}\|\int_{t_{j}}^{t_{j+1}}(\mathbb{E}\|A^{(\beta-1)/2}f(u(\tau))\|^{2})^{1/2}d\tau\\ &\leq Ck\sum_{j=0}^{n-1}\|F_{n-j}A^{(1-\beta)/2}\|\\ &\leq C\sqrt{n}k\left(\sum_{j=0}^{n-1}\|F_{j}A^{(1-\beta)/2}\|^{2}\right)^{1/2}\end{split}$
@@
$\begin{split}Err_{d,3}^{2}&\leq Ck\sum_{j=0}^{n-1}\|F_{j}A^{(1-\beta)/2}\|^{2}\\ &\leq Ck\sum_{j=0}^{n-1}\left(\sup_{v\neq 0}\frac{\|F_{j}A^{(1-\beta)/2}v\|}{\|v\|}\right)^{2}\\ &=C\sup_{v\neq 0}\frac{k\sum_{j=0}^{n-1}\|F_{j}A^{(1-\beta)/2}v\|^{2}}{\|v\|^{2}}\\ &\leq C\sup_{v\neq 0}\frac{(k^{\beta}+h^{2\beta})\|A^{(1-\beta)/2}v\|^{2}_{\beta-1}}{\|v\|^{2}}\\ &\leq C(k^{\beta}+h^{2\beta}).\end{split}$
@@
$\begin{split}Err_{d,4}&=\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}(\mathbb{E}\|(E(t_{n}-t_{j})-E(t_{n}-\tau))A^{(1-\beta)/2}A^{(\beta-1)/2}f(u(\tau))\|^{2})^{1/2}d\tau\\ &\leq C\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}\|(E(t_{n}-t_{j})-E(t_{n}-\tau)))A^{(1-\beta)/2}\|(\mathbb{E}\|A^{(\beta-1)/2}f(u(\tau))\|^{2})^{1/2}d\tau\\ &\leq C\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}\|(E(t_{n}-t_{j})-E(t_{n}-\tau))A^{(1-\beta)/2}\|d\tau\\ &\leq C\sqrt{n}\left(\sum_{j=0}^{n-1}\left(\int_{t_{j}}^{t_{j+1}}\|(E(t_{n}-t_{j})-E(t_{n}-\tau))A^{(1-\beta)/2}\|d\tau\right)^{2}\right)^{1/2}\\ &\leq C\sqrt{n}\sqrt{k}\left(\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}\|(E(t_{n}-t_{j})-E(t_{n}-\tau))A^{(1-\beta)/2}\|^{2}d\tau\right)^{1/2}\end{split}$
@@
$\begin{split}&=2\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}\|(R(kA_{h})^{n-j}\mathcal{P}_{h}-E(t_{n}-t_{j}))A^{(1-\beta)/2}A^{(\beta-1)/2}Q^{1/2}\|_{HS}^{2}d\tau\\ &+2\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}\|(E(t_{n}-t_{j})-E(t_{n}-\tau))A^{(1-\beta)/2}A^{(\beta-1)/2}Q^{1/2}\|_{HS}^{2}d\tau\\ &\leq Ck\sum_{j=0}^{n-1}\|F_{j}A^{(1-\beta)/2}\|^{2}+2\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}\|(E(t_{n}-t_{j})-E(t_{n}-\tau))A^{(1-\beta)/2}\|^{2}d\tau\\ &\leq C(k^{\beta}+h^{2\beta})+Ck^{\beta}\leq C(k^{\beta}+h^{2\beta}).\end{split}$
@@
$\mathbb{E}\|e^{n}_{v}\|^{2}\leq C(k^{\gamma}+h^{2\beta})+Ck\sum_{J=0}^{n-1}\mathbb{E}\|U^{j}-u(t_{j})\|^{2}.$ (4.7)
@@
$\begin{split}\mathbb{E}\|U^{n}-u(t_{n})\|^{2}&=\mathbb{E}\|(\mathcal{P}_{h}+A_{h})^{-1}V^{n}-(I+A)^{-1}v(t_{n})\|^{2}\\ &=\mathbb{E}\|(\mathcal{P}_{h}+A_{h})^{-1}V^{n}-(I+A)^{-1}v(t_{n})\|^{2}\\ &=\mathbb{E}\|(\mathcal{P}_{h}+A_{h})^{-1}V^{n}-(I+A)^{-1}V^{n}+(I+A)^{-1}V^{n}-(I+A)^{-1}v(t_{n})\|^{2}\\ &\leq 2\mathbb{E}\|(\mathcal{P}_{h}+A_{h})^{-1}V^{n}-(I+A)^{-1}V^{n}\|^{2}+\mathbb{E}\|V^{n}-v(t_{n})\|^{2}\end{split}$
By using the connection between the resolvent of the generator and the corresponding $C_{0}$-semigroup [13, Theorem 4.2], for $n\geq 0$, we get
@@
$\begin{split}\mathbb{E}\|U^{n}-u(t_{n})\|^{2}&\leq C(k^{\gamma}+h^{2\beta})+Ck\sum_{J=0}^{n-1}\mathbb{E}\|U^{j}-u(t_{j})\|^{2}\\ &\leq C(k^{\gamma}+h^{2\beta})e^{Ckn}\\ &\leq C(k^{\gamma}+h^{2\beta})\end{split}$
@@ 5. Numerical Experiments
We consider the following stochastic semilinear generalized Benjamin-Bona-Mahony (BBM) equation
$\begin{cases}du-d(\Delta u)-(\Delta u-\Delta^{2}u)dt=udt+dW\quad\text{ in }(0,1)\times(0,1),\\ u(t,0)=u(t,1)=0\hskip 142.26378ptt\in(0,1),\\ \Delta u(t,0)=\Delta u(t,1)=0\hskip 130.88284ptt\in(0,1),\\ u(0,x)=\sin(\frac{2\pi x}{1+4\pi^{2}})\hskip 150.79968ptx\in(0,1).\end{cases}$ (5.1)
@@
$\begin{cases}dv-\Delta vdt=udt+dW\quad\text{ in }(0,1)\times(0,1),\\ v=u-\Delta u\hskip 88.2037pt\text{ in }(0,1)\times(0,1),\\ v(t,0)=u(t,0)=0\hskip 56.9055ptt\in(0,1),\\ v(0,x)=\sin(2\pi x)\hskip 65.44142ptx\in(0,1).\end{cases}$ (5.2)
@@
Using the full discretization scheme, we perform numerical computation for the example (5.2) to support Theorem. 2.6. Below, we present Figures 2 and 2, which describe the rate of convergence for the equation (5.2).
@@
    The rate of strong convergence of finite element approximations of stochastic semilinear generalized BBM equation with respect to the space-discretized parameter $h:$ We take $\beta=1,\,d=1$, hence $s>1-\frac{1}{2}$. We choose $s=1/2+.0005$, $h_{ref}=1024$ (step length of space discretization for reference solution) and $k=.01$ k =0.01 (time step to be fixed) and $10^{3}$ Monte-Carlo samples for sampling in the stochastic case; see Figure 2.
@@
    The rate of strong convergence of time-discretized approximations of stochastic semilinear generalized BBM equations with respect to the time-discretized parameter $k$: We take $\beta=0.5$,$d=1$, hence $s>0$. We choose $s=0.0005$, $h_{ref}=2^{-6}$ (step length to be fixed ) and $k_{ref}=2^{-16}$(step length of time discretization for reference solution) and $10^{2}$ Monte-Carlo samples for sampling in the stochastic case; Figure 2.

### v2

@@ 4. Proof of Main Results
$\begin{split}\mathbb{E}\|u_{h}(t)\|^{2}&=\mathbb{E}\|(\mathcal{P}_{h}+A_{h})^{-1}v_{h}(t)\|^{2}\\ &\leq\mathbb{E}\|v_{h}(t)\|^{2}\\ &\leq B(t)+3Ct\int_{0}^{t}\mathbb{E}\|u_{h}(\tau)\|^{2}d\tau.\end{split}$
@@
Now we prove our first main result of the paper. First, we find the error between $v(t)$ and $v_{h}(t)$. Using this error, we obtain the rate of convergence of the finite element approximation of the fourth-order stochastic pseudo-parabolic equation (2.1).
@@
(4.3) $I_{1}(t)\leq Ch^{\beta}\left(\|v_{0}\|_{L_{2}(\Omega,\dot{H}^{\beta})}+\|A^{(\beta-1)/2}Q^{1/2}\|\right).$
@@
[…]
$\begin{split}&\leq Ch^{\beta}\int_{0}^{t}(t-\tau)^{-\beta/2}(1+\|u(\tau)\|_{L_{2}(\Omega,H)})d\tau\\ &+C\int_{0}^{t}\|u(\tau)-u_{h}(\tau)\|_{L_{2}(\Omega,H)}d\tau\\ &\leq Ch^{\beta}+C\int_{0}^{t}\|u(\tau)-u_{h}(\tau)\|_{L_{2}(\Omega,H)}d\tau.\end{split}$
@@
To estimate $J(t)$, we use the connection between the resolvent of the generator and the corresponding $C_{0}$-semigroup [26, Chapter 3 Theorem 4.2]. For $t\geq 0$, we get
@@
This completes the proof of Theorem 2.3. ∎
@@
The solution of the stochastic parabolic component $v(t)$ is regular in time ( see [37, Lemma 2.7] and [28, Proposition 3.4]) i.e.
@@
[…]
$\begin{split}&=K_{f}Ck\sum_{j=0}^{n-1}\left(\mathbb{E}\|U^{j}\|^{2}\right)^{1/2}\\ &\leq K_{f}Ck\sqrt{n}\left(\sum_{j=0}^{n-1}\mathbb{E}\|U^{j}\|^{2}\right)^{1/2}.\end{split}$
@@
(4.5) $\|A^{\nu}E(t)\|\leq Ct^{-\nu},\quad\text{ for }t>0,$
(4.6) $\|A^{\mu}(I-E(t))\|\leq Ct^{\mu},\quad\text{ for }t\geq 0.$
Now we prove our second main result. First, we will compute the full discretization error between $V^{n}$ and $v(t_{n})$ for all $t_{n}\in[0,T]$. By using this error, we will find the full discretization error for the fourth-order stochastic pseudo-parabolic equation.
@@
[…]
$\begin{split}&\leq CK_{f}k\sum_{j=0}^{n-1}(\mathbb{E}\|U^{j}-u(t_{j})\|^{2})^{1/2}\\ &\leq C\sqrt{n}k\left(\sum_{j=0}^{n-1}\mathbb{E}\|U^{j}-u(t_{j})\|^{2}\right)^{1/2}.\end{split}$
@@
$\begin{split}Err_{d,2}&=\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}(\mathbb{E}\|R(kA_{h})^{n-j}\mathcal{P}_{h}(f(u(t_{j}))-f(u(\tau)))\|^{2})^{1/2}d\tau\\ &\leq CK_{f}\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}(\mathbb{E}\|u(t_{j})-u(\tau)\|^{2})^{1/2}d\tau\\ &\leq CK_{f}\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}(t_{j}-\tau)^{\gamma/2}d\tau\\ &\leq Ck^{\gamma/2}.\end{split}$
For the estimate of $Err_{d,3}$, we apply Lemma $2.8$ of [37] and linear growth condition (ii) of the function $f$ to obtain
$\begin{split}Err_{d,3}&=\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}(\mathbb{E}\|R(kA_{h})^{n-j}\mathcal{P}_{h}-E(t_{n}-t_{j})f(u(\tau))\|^{2})^{1/2}d\tau\\ &=\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}(\mathbb{E}\|F_{n-j}A^{(1-\beta)/2}A^{(\beta-1)/2}f(u(\tau))\|^{2})^{1/2}d\tau\\ &\leq\sum_{j=0}^{n-1}\|F_{n-j}A^{(1-\beta)/2}\|\int_{t_{j}}^{t_{j+1}}(\mathbb{E}\|A^{(\beta-1)/2}f(u(\tau))\|^{2})^{1/2}d\tau\\ &\leq Ck\sum_{j=0}^{n-1}\|F_{n-j}A^{(1-\beta)/2}\|\\ &\leq C\sqrt{n}k\left(\sum_{j=0}^{n-1}\|F_{j}A^{(1-\beta)/2}\|^{2}\right)^{1/2}.\end{split}$
@@
[…]
$\begin{split}&=C\sup_{v\neq 0}\frac{k\sum_{j=0}^{n-1}\|F_{j}A^{(1-\beta)/2}v\|^{2}}{\|v\|^{2}}\\ &\leq C\sup_{v\neq 0}\frac{(k^{\beta}+h^{2\beta})\|A^{(1-\beta)/2}v\|^{2}_{\beta-1}}{\|v\|^{2}}\\ &\leq C(k^{\beta}+h^{2\beta}).\end{split}$
@@
$\begin{split}Err_{d,4}&=\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}(\mathbb{E}\|(E(t_{n}-t_{j})-E(t_{n}-\tau))A^{(1-\beta)/2}A^{(\beta-1)/2}f(u(\tau))\|^{2})^{1/2}d\tau\\ &\leq C\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}\|(E(t_{n}-t_{j})-E(t_{n}-\tau)))A^{(1-\beta)/2}\|(\mathbb{E}\|A^{(\beta-1)/2}f(u(\tau))\|^{2})^{1/2}d\tau\\ &\leq C\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}\|(E(t_{n}-t_{j})-E(t_{n}-\tau))A^{(1-\beta)/2}\|d\tau\\ &\leq C\sqrt{n}\left(\sum_{j=0}^{n-1}\left(\int_{t_{j}}^{t_{j+1}}\|(E(t_{n}-t_{j})-E(t_{n}-\tau))A^{(1-\beta)/2}\|d\tau\right)^{2}\right)^{1/2}\\ &\leq C\sqrt{n}\sqrt{k}\left(\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}\|(E(t_{n}-t_{j})-E(t_{n}-\tau))A^{(1-\beta)/2}\|^{2}d\tau\right)^{1/2}.\end{split}$
@@
$\begin{split}&\leq\sum_{j=0}^{n-1}2\int_{t_{j}}^{t_{j+1}}\|(R(kA_{h})^{n-j}\mathcal{P}_{h}-E(t_{n}-t_{j}))Q^{1/2}\|_{HS}^{2}d\tau\\ &+\sum_{j=0}^{n-1}2\int_{t_{j}}^{t_{j+1}}\|(E(t_{n}-t_{j})-E(t_{n}-\tau))Q^{1/2}\|_{HS}^{2}d\tau\\ &=2\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}\|(R(kA_{h})^{n-j}\mathcal{P}_{h}-E(t_{n}-t_{j}))A^{(1-\beta)/2}A^{(\beta-1)/2}Q^{1/2}\|_{HS}^{2}d\tau\\ &+2\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}\|(E(t_{n}-t_{j})-E(t_{n}-\tau))A^{(1-\beta)/2}A^{(\beta-1)/2}Q^{1/2}\|_{HS}^{2}d\tau\\ &\leq Ck\sum_{j=0}^{n-1}\|F_{j}A^{(1-\beta)/2}\|^{2}+2\sum_{j=0}^{n-1}\int_{t_{j}}^{t_{j+1}}\|(E(t_{n}-t_{j})-E(t_{n}-\tau))A^{(1-\beta)/2}\|^{2}d\tau\\ &\leq C(k^{\beta}+h^{2\beta})+Ck^{\beta}\leq C(k^{\beta}+h^{2\beta}).\end{split}$
@@
(4.7) $\mathbb{E}\|e^{n}_{v}\|^{2}\leq C(k^{\gamma}+h^{2\beta})+Ck\sum_{J=0}^{n-1}\mathbb{E}\|U^{j}-u(t_{j})\|^{2}.$
@@
$\begin{split}\mathbb{E}\|U^{n}-u(t_{n})\|^{2}&=\mathbb{E}\|(\mathcal{P}_{h}+A_{h})^{-1}V^{n}-(I+A)^{-1}v(t_{n})\|^{2}\\ &=\mathbb{E}\|(\mathcal{P}_{h}+A_{h})^{-1}V^{n}-(I+A)^{-1}v(t_{n})\|^{2}\\ &=\mathbb{E}\|(\mathcal{P}_{h}+A_{h})^{-1}V^{n}-(I+A)^{-1}V^{n}+(I+A)^{-1}V^{n}-(I+A)^{-1}v(t_{n})\|^{2}\\ &\leq 2\mathbb{E}\|(\mathcal{P}_{h}+A_{h})^{-1}V^{n}-(I+A)^{-1}V^{n}\|^{2}+\mathbb{E}\|V^{n}-v(t_{n})\|^{2}.\end{split}$
By using the connection between the resolvent of the generator and the corresponding $C_{0}$-semigroup [26, Chapter 3 Theorem 4.2], for $n\geq 0$, we get
@@
$\begin{split}\mathbb{E}\|U^{n}-u(t_{n})\|^{2}&\leq C(k^{\gamma}+h^{2\beta})+Ck\sum_{J=0}^{n-1}\mathbb{E}\|U^{j}-u(t_{j})\|^{2}\\ &\leq C(k^{\gamma}+h^{2\beta})e^{Ckn}\\ &\leq C(k^{\gamma}+h^{2\beta}).\end{split}$
@@ 5. Numerical Experiments
We consider the following fourth-order stochastic pseudo-parabolic equation
(5.1) $\begin{cases}du-d(\Delta u)-(\Delta u-\Delta^{2}u)dt=udt+dW\quad\text{ in }(0,1)\times(0,1),\\ u(t,0)=u(t,1)=0\hskip 136.5733ptt\in(0,1),\\ \Delta u(t,0)=\Delta u(t,1)=0\hskip 113.81102ptt\in(0,1),\\ u(0,x)=\frac{1}{1+4\pi^{2}}\sin(2\pi x)\hskip 113.81102ptx\in(0,1).\end{cases}$
@@
(5.2) $\begin{cases}dv-\Delta vdt=udt+dW\quad\text{ in }(0,1)\times(0,1),\\ v=u-\Delta u\hskip 88.2037pt\text{ in }(0,1)\times(0,1),\\ v(t,0)=u(t,0)=0\hskip 56.9055ptt\in(0,1),\\ v(0,x)=\sin(2\pi x)\hskip 65.44142ptx\in(0,1).\end{cases}$
@@
Using the full discretization scheme, we perform numerical computation for the example (5.2) to support Theorem. 2.6. Below, we present FIGURE 2 and 2, which describe the rate of convergence for the equation (5.2).
@@
    The rate of strong convergence of finite element approximations of fourth-order stochastic pseudo-parabolic equation with respect to the space-discretized parameter $h:$ We take $\beta=1,\,\,d=1$, hence $s>1-\frac{1}{2}$. We choose $s=1/2+.0005$, $h_{ref}=1024$ (step length of space discretization for reference solution) and $k=.01$ k =0.01 (time step to be fixed) and $10^{3}$ Monte-Carlo samples for sampling; see FIGURE 2.
@@
    The rate of strong convergence of time-discretized approximations of fourth-order stochastic pseudo-parabolic equations with respect to the time-discretized parameter $k$: We take $\gamma=0.98$, $d=1$, hence $s>0$. We choose $s=1/2+0.0005$, $h_{ref}=2^{-5}$ (step length to be fixed ) and $k_{ref}=2^{-17}$(step length of time discretization for reference solution) and $10^{2}$ Monte-Carlo samples for sampling; FIGURE 2.